WAVE INTERFERENCE: Discuss interference, constructive and destructive interference. List the phase and path difference equations for constructive and destructive conditions. Discuss the phase changes condition due to reflection of light from a surface. Summarize equations of interference for thin film. Give real life examples for interference
Answer #1

t is phenomenon - redistribution interference of light of light energy in the medium on the account of two Cho coherent sourcY= R sin (wtto) Resultant wave, with Ramplitude Resultant Ware intensity - R² = R² - Asintotto) R² (6530 + sinho) a3 +6² 632dPath difference X= (2n-1) ITXA = 2 TT Il (2n-1) a Where n = 1, 2, 3 2 reflected by Real life examples light this oil film flo